OpenStudy (anonymous):

sam rolls a fair 4-sided die containing the numbers 1,2,3, and 4. Tyler rolls a fair 6-sided die containing the numbers 1,2,3,4,5 and 6. What is the probability that sam rolls a larger number than Tyler?

OpenStudy (anonymous):

1/4 i suppose if i am not wrong

OpenStudy (anonymous):

thanks :) but, how did you get it?

OpenStudy (anonymous):

see we hav favourable items as (2,1) (3,1) (3,2) (4,1) (4,2) (4,3) and total cases are - 4 * 6 =24 thus prob of sam rolls higher = 6/24 =1/4

OpenStudy (anonymous):

ohhhh thanks!! i get it now

OpenStudy (anonymous):

you're welcome